OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (5, -5) with 0° ≤ θ < 360°. (5 √2, 315°), (-5 √2, 135°) (5 √2, 45°), (-5 √2, 225°) (5 √2, 135°), (-5 √2, 315°) (5 √2, 225°), (-5 √2, 45°)

OpenStudy (anonymous):

x= rcos(theta) y= rsin(theta) arctan(y/x)= theta

OpenStudy (anonymous):

What does that mean? Please explain

OpenStudy (anonymous):

Ok, do you how to convert from cartesian to polar coordinates?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Ok, so how do you go from x= 5 to r= ?

OpenStudy (anonymous):

um, you would use x=rcos(theta) to suggest that 5=rcos(theta) then get r by itself?

OpenStudy (anonymous):

Umm, 5 = rcos(theta)

OpenStudy (anonymous):

x^2 + y^2 = r^2 thus you can find your r value and theta

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

The question is asking you to find r and theta

OpenStudy (anonymous):

Ill walk you through it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):